OpenStudy (anonymous):

Find the exact value sin^1 0

OpenStudy (anonymous):

\[\sin^{-1}(0)=0\]

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

Yeah, its 0. Because you want "the angle that makes sin(x) zero. Which is zero.

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

wat bout this one

OpenStudy (anonymous):

You want the "angle that makes sin sqrt(2)/2". You know that sin(pi/4)=sqrt(2)/2. so the ARCSIN(sqrt(2)/2) is pi/4.

OpenStudy (anonymous):

pi/4 You should really memorize the values of the special trig functions -- they pop up everywhere and you can derive others from them quite easily when you're taking a test http://www.analyzemath.com/trigonometry/special_angles.html

OpenStudy (anonymous):

thank u guys

OpenStudy (anonymous):

wat abut cos

OpenStudy (anonymous):

pi/3 http://www.analyzemath.com/trigonometry/special_angles.html

OpenStudy (anonymous):